OpenStudy (anonymous):

1+(1/x) / 1-(1/x) = ?/?

OpenStudy (anonymous):

multiply top and bottom by x

OpenStudy (anonymous):

\[\frac{1+\frac{1}{x}}{1-\frac{1}{x}}\times \frac{x}{x}\]

OpenStudy (anonymous):

gives \[\frac{x+1}{x-1}\] just like that

OpenStudy (anonymous):

k?